in a given ellipse, the distance between foci is 2c and between vertices 2a. determine the distance from one of the foci to the nearest vertex
Answer by Edwin McCravy(20081)




The vertices are P and S
The foci are Q and R

Given: QR = 2c
       PS = 2a

To find: PQ = RS = x

     PS = PQ + QR + RS

     2a =  x + 2c +  x

     2a = 2x + 2c

2a - 2c = 2x

  a - c = x

Answer: a - c

Edwin